the current i is passing through a conductor of length L . if S is the specific charge of electron then the total ,momentum of electrons in conductors is. 
​ 1) iL/S
​2)  2iL/S
​3)  iS/L
​4) iL/S

Grade:12th pass

1 Answers

Khimraj
5 years ago
Acoording to question q/m = S and q/t = i and v = L/t
Then momentum p = m*v  = (q/S)*(L/t). Putting value of t
p = (q/S)*(L*i/q) = iL/S.
Hope it clears.